OpenStudy (anonymous):

A rabbit and a turtle were to start a race at the same place. The rabbit was asleep when the turtle started. He slept for seven hours before he started the race. The rabbit's average rate when he was awake was 5 km/hr. The turtle never stopped for a break, but his pace was only 1km/hr. How long will the rabbit run before it catches the turtle?

OpenStudy (anonymous):

how many hours?

OpenStudy (anonymous):

Let t be the time for the rabbit, r, to catch the turtle, t. The turtle has a head start of 7*1km/hr. r t = d 5 t = 7 + t\[t=\frac{7}{4}\text{ hours} \]

OpenStudy (anonymous):

thanks robtobey

OpenStudy (anonymous):

Your are welcome., Thank you for the medal(s).

OpenStudy (anonymous):

As an aside, the equation should have been written as follows although the end result would have been the same because of the turtle's rate of 1km/hr. 5 t=7+1t